# Sharp Thresholds for Factors in Random Graphs

### v1

@@ 1 Introduction
What is the simplest, local obstruction for a graph $G$, on an even number of vertices, to admit a perfect matching? Well, if an isolated vertex exists, a perfect matching cannot. Back in 1966, Erdős and Rényi [9] showed that this simplest, local obstruction is indeed the only relevant obstruction, when it comes to the random graph $G(n,p)$, with vertices $[n]=\{1,\dots,n\}$, and where each edge is included independently with probability $p$. To be precise, recall that a sequence $p^{*}=p^{*}_{n}$ is a sharp threshold for a graph property $\mathcal{E}=\mathcal{E}_{n}$ if for all $\varepsilon>0$ and $p\leq(1-\varepsilon)p^{*}$ the random graph $G(n,p)$ does not have $\mathcal{E}$ with high probability (whp), that is, with probability tending to one as $n$ increases, while for all $p\geq(1+\varepsilon)p^{*}$ the random graph $G(n,p)$ does have $\mathcal{E}$ whp. Erdős and Rényi [9] observed that the sharp threshold for the existence of perfect matchings coincides with the sharp threshold for the disappearance of isolated vertices.
@@
This is an application of Theorem 3.22 (i) in [15]. ∎
@@ 1.1 Main Result
$\pi^{\prime}=1-(1-\pi)^{r!/\operatorname{aut}(F)}.$
@@ 2.1 Induced Edges, Avoidable Configurations And Clean Cycles
When trying to construct a coupling between the random $F$-graph $H={H_{F}}(n,\pi)$ and the copies of $F$ in the random graph $G=G(n,p)$, one invariably encounters the following fundamental problem. Let $H^{\prime}$ be the $F$-graph of $G$, meaning that we include each copy of $F$ in $G$ as an $F$-edge in $H^{\prime}$. For all $10<\pi<$ 1, $H$ can clearly attain any $F$-graph. But this is not the case for $H^{\prime}$, due to the existence of so-called ’induced $F$-edges’: Assume that there is an $F$-edge $h$, meaning a copy $h$ of $F$, which is not contained in $H$, but that for each (graph) edge $e$ in the copy $h$ of $F$ there exists an $F$-edge $h_{e}$ in $H$ that covers $e$. Then $H^{\prime}$ cannot attain $H$, because if the copies $h_{e}$ of $F$ appear in $G$, then so does the copy $h$ of $F$.
@@ 2.2 Sparse Clean Cycles And Dummy Edges
This leaves us with clean $F$-cycles, which may exist in $H$. Riordan avoids the discussion of these clean $F$-cycles by restricting to the class of graphs $F$ from Theorem 1.3, which he called ‘nice’: A graph with these properties cannot be induced by a clean $F$-cycle. Heckel [13], on the other hand, tweaks Riordan’s coupling to deal with the only type of clean $K_{3}$-cycle which can induce a triangle $K_{3}$, namely the one of length $3$. Since we consider the case of general strictly 1-balanced graphs, we also have to deal with clean $F$-cycles. But as opposed to Heckel, we have to consider all types of clean $F$-cycles. While this is a major complication in its own right, there is one particular type of clean $F$-cycle which is particularly demanding: so-called ‘sparse clean $F$-cycles’, which are given by two $F$-edges, meaning two copies of $F$, that share exactly one (graph) edge (and no other vertices).
@@ 2.3 Coupling Clean Cycles
With this modification in place, we turn back to the proof in [13], where Heckel couples the clean $F$-cycles in $H$ and their shadow graphs in $G$ so that they match up whp. As opposed to Heckel, we have to couple all types of clean $F$-cycles in $H$, not just one type, with their (selected) shadow graphs in $G^{*}$. For this purpose, as in [13], we have to control the overlaps of pairs of clean $F$-cycles, and the overlaps of their (selected!) shadow graphs in $G^{*}$. Heckel used an exhaustive case distinction and direct computation to gain control. This is not an option for the general case, which brings us to our second main contribution. For the discussion of two overlapping, distinct clean $F$-cycles as well as the discussion of other more complex, composed $F$-graphs, we observe that these are in fact avoidable configurations. Thus, it is Riordan’s work on avoidable configurations that immediately gives us control over the overlaps. On the graph side, we establish that shadow graphs $S$ of clean $F$-cycles are strictly balanced, that is, the edge density $e(S)/v(S)$ of any proper subgraph is strictly smaller than the edge density of $S$. Here, it is yet again crucial that we consider the dummy edge when we discuss the shadow graphs of sparse clean $F$-cycles. With this property in place, we easily control the overlaps in the graph case. As with the avoidable configurations above, the fact that these shadow graphs are strictly balanced also helps later on in the proof.
@@ 3.1 Global Constants
3.2 A New Type of Hypergraph
An $F$-graph $H=(V,E)$ is given by the vertex set $V$ and a set $E$ of $F$-edges, that is, copies of $F$ with vertices in $V$. The shadow graph $G=(V^{\prime},E^{\prime})$ of $H$ is the graph with vertices $V^{\prime}=V$ and the edge set $E^{\prime}$, which is the union of the edge sets of all $F$-edges in $H$. Formally, let $h=(V_{h},E_{h})$ be the $F$-edges $h\in E$, then we have
@@ 3.2 A New Type of Hypergraph
For a graph $G=(V,E)$, let the $F$-graph $H=(V^{\prime},E^{\prime})$ of $G$ be given by the vertices $V^{\prime}=V$ and the set $E^{\prime}$ of copies of $F$ in $G$. Finally, note that we may think of $r$-uniform hypergraphs as $K_{r}$-graphs and of simple graphs as $K_{2}$-graphs. Thus, it suffices to introduce notions for $F$-graphs (and we take the liberty to allow $F=K_{2}$ for definitions).
@@ 3.3 Induced Edges, Avoidable Configurations And Clean cycles
$n(H)=(r-1)e(H)+c(H)-v(H).$
We will tacitly iterate the following observations on several occasions, to identify minimizers of the nullity in classes of $F$-graphs. Consider two connected $F$-graphs $H_{1}$, $H_{2}$. Let $v$ be a vertex in their vertex overlap such that no $F$-edge incident to $v$ is contained in both $H_{1}$ and $H_{2}$. Obtain the copy $H^{\prime}_{2}$ of $H_{2}$ by substituting $v$ with an entirely new vertex. Then we have $n(H_{1}\cup H^{\prime}_{2})=n(H_{1}\cup H_{2})$ if $v$ is the only vertex in the overlap and $n(H_{1}\cup H^{\prime}_{2})=n(H_{1}\cup H_{2})-1$ otherwise.
Now, let $h$ be an $F$-edge in the overlap of $H_{1}$ and $H_{2}$ and $v$ a vertex in $h$ such that each of the remaining $(r-1)$ vertices in $h$ is not incident to any other $F$-edge in the overlap. Obtain the copy $H^{\prime}_{2}$ of $H_{2}$ by substituting these $(r-1)$ vertices with entirely new vertices. Then we have $n(H_{1}\cup H^{\prime}_{2})=n(H_{1}\cup H_{2})$.
Next, let an $F$-edge $h$ be induced by $H$ if it is not contained in $H$, but it is contained in the $F$-graph of the shadow graph of $H$. An $F$-graph $H$ is an avoidable configuration if it is connected and $n(H)\geq 2$. Further, for $k>2$, the $F$-graph $H$ is a clean $F$-cycle of length $k$ if the following holds. There exists an ordering $h_{1},\dots,h_{k}$ of the $F$-edges of $H$ and there are distinct vertices $v_{1},\dots,v_{k}$ such that $h_{i}$ and $h_{i+1}$ overlap exactly in $v_{i}$ for all $i\leq k$ (using $h_{k+1}=h_{1}$), and $h_{i}$, $h_{j}$ do not overlap otherwise. For $k=2$, the $F$-graph $H$ is a clean $F$-cycle of length $2$ if it has exactly two $F$-edges, which overlap in exactly two vertices. The next result from [24] shows that we can control induced $F$-edges by controlling avoidable configurations and clean $F$-cycles. [^1] Note that the statement in [24] is slightly different: the conclusion is only that there exists an avoidable configuration in $H$, or a clean $F$-cycle that induces $E$. However, when following through the proof in [24] it is clear that the avoidable configuration that is found indeed induces $E$.
@@
3.4 A New Type of Graph
Let $C$ be a clean $F$-cycle. We say that $C$ is sparse if $e(C)=2$ and the two overlap vertices $v_{1},v_{2}$ form an edge $e=\{v_{1},v_{2}\}$ in both $F$-edges $h_{1},h_{2}$ of $C$, that is, the edge $e$ is contained in the copy $h_{1}$ of $F$ as well as in the copy $h_{2}$ of $F$. Otherwise, we say that the clean $F$-cycle $C$ is dense.
@@ 3.4 A New Type of Graph
In order to match ${\mathbb{E}}[X]$ and ${\mathbb{E}}[Y]$, we select only a fraction $p$ of the sparse clean $F$-cycles in $G(n,p)$. To model this selection, we introduce the following ‘dummy edges’. A d-graph $G=(V,E)$ is given by vertices $V$ and an edge set $E=U\cup D$, where the (usual) edges $U$ are some subsets of $V$ of size $2$, and the dummy edges $D$ are some sparse clean $F$-cycles with vertices in $V$. The d-graph $G$ is a clean d-cycle if it is the shadow graph of a dense clean $F$-cycle or if it consists of all the edges of the shadow graph $S$ of a sparse clean $F$-cycle $C$, and, additionally, the dummy edge corresponding to $C$. This definition ensures that for any clean d-cycle $C$ of length $k$ we have $ke(F)$ edges. Finally, the random d-graph $G^{*}(n,p)$ has the vertices $[n]$, and each edge (including dummy edges) is included independently with probability $p$. Note that we obtain $G(n,p)$ from $G^{*}(n,p)$ by discarding the dummy edges.
For d-graphs $G$ we also use $v(G),e(G),c(G)$ as before, where $G$ is connected if $G$ without the dummy edges is. We also use the 1-density $d_{1}(G)$ from Equation (1) and the edge density $e(G)/v(G)$, which means that also strictly 1-balanced and strictly balanced d-graphs are defined. The next result is a key argument in the proof.
@@
Now, assume that $G$ is not strictly balanced and let $S\subsetneqq G$ be strictly balanced with maximal density. Assume that $S$ does not cover one of the overlap vertices $v_{i}$. Then we can split at $v_{i}$ to see that $S$ is a subgraph of the shadow graph of a clean $F$-path $P$ with copies $F^{\prime}_{1},\dots,F^{\prime}_{\ell}$ of $F$, that is, $F^{\prime}_{i}$ and $F^{\prime}_{i+1}$ overlap in exactly one vertex for $i<\ell$ and $F^{\prime}_{i}$, $F^{\prime}_{j}$ have no overlap otherwise. Since $S$ is connected, we may assume that $S$ covers all overlap vertices of $P$ (possibly $\ell=1$), and that, using $S_{i}=F^{\prime}_{i}\cap S$, we have $v(S_{i})\geq 2$ for all $i$. It follows that
@@ 4.1 Overlaps
Note that $C_{1}\cup C_{2}$ is an avoidable configuration, so Lemma 3.2 completes the proof. ∎
@@
Let $C_{i}$ be the clean $F$-cycle corresponding to $G_{i}$. Further, let $S=G_{1}\cap G_{2}$ be the intersection graph, then we have $\mathbb{E}[X]=O(n^{f(S)})$ with
@@ 4.2 Coupling
$4\left(\sum_{i}\sum_{j\in B_{i}}\mathbb{E}[X_{i}]\mathbb{E}[X_{j}]+\sum_{i}\sum_{j\in B_{i}\setminus\{i\}}\mathbb{E}[X_{i}X_{j}]\right).$
@@
For a clean $F$-cycle $C$ let $X_{C}$ be the indicator for the event that $C$ is in ${H_{F}}(n,\pi)$. Further, let $B_{C}$ be the set of all clean $F$-cycles that are not vertex disjoint from $C$. Note that for $C^{\prime}\in B_{C}$ we have $\mathbb{E}[X_{C}X_{C^{\prime}}]\geq\mathbb{E}[X_{C}]\mathbb{E}[X_{C^{\prime}}]$, so Lemma 4.1, as well as Lemma 3.2 and the fact that $n(C)=1$ (to handle the case $C^{\prime}=C$), together with Theorem 4.3 yield a coupling of $(X_{C})_{C}$ and independent Poisson variables $(Y_{C})_{C}$ with parameters $\mathbb{E}[X_{C}]$ such that $(X_{C})_{C}=(Y_{C})_{C}$ whp. We proceed analogously for $G^{*}(n,p)$ to arrive at a coupling of the indicators $(X^{\prime}_{C})_{C}$ and independent Poisson variables $(Y^{\prime}_{C})_{C}$ with parameters $\mathbb{E}[X^{\prime}_{C}]$ such that $(X^{\prime}_{C})_{C}=(Y^{\prime}_{C})_{C}$ whp. Now, note that $\mathbb{E}[X_{C}]=(1-n^{-\delta})^{e(C)}\mathbb{E}[X^{\prime}_{C}]$ and that $\mathbb{E}[|\mathcal{C}_{i}|]=O(n^{e(F)\varepsilon})$, so we can couple $(Y_{C})_{C}$ with $(Y^{\prime}_{C})_{C}$ such that $(Y_{C})_{C}=(Y^{\prime}_{C})_{C}$ whp. ∎

### v2

@@ 1 Introduction
What is the simplest, local obstruction for a graph $G$, on an even number of vertices, to admit a perfect matching? Well, if an isolated vertex exists, a perfect matching cannot. Back in 1966, Erdős and Rényi [9] showed that this simplest, local obstruction is indeed the only relevant obstruction, when it comes to the random graph $G(n,p)$, with vertices $n]=\{1,\dots,n\}$, and where each edge is included independently with probability $p$. To be precise, recall that a sequence $p^{*}=p^{*}_{n}$ is a sharp threshold for a graph property $\mathcal{E}=\mathcal{E}_{n}$ if for all $\varepsilon>0$ and $p\leq(1-\varepsilon)p^{*}$ the random graph $G(n,p)$ does not have $\mathcal{E}$ with high probability (whp), that is, with probability tending to one as $n$ increases, while for all $p\geq(1+\varepsilon)p^{*}$ the random graph $G(n,p)$ does have $\mathcal{E}$ whp. Erdős and Rényi [9] observed that the sharp threshold for the existence of perfect matchings coincides with the sharp threshold for the disappearance of isolated vertices.
@@
This is an application of Theorem 3.22 (i) in [15], originally due to [27]. ∎
@@ 1.1 Main Result
$\pi^{\prime}=1-(1-\pi)^{r!/\operatorname{aut}(F)}.$ (2)
Remark 1.7.
It follows from the proof that Theorems 1.5 and 1.6 remain valid if $F$ is a strictly 1-balanced $u$-uniform hypergraph. In this case, $G(n,p)$ in these theorems needs to be replaced by $H_{u}(n,p)$, the random $u$-uniform hypergraph. In fact, the proof for $u>2$ can be simplified, since two copies of $F$ overlapping in exactly two vertices can never intersect in an $u$-edge. Thus, the obstacle posed by sparse clean cycles as defined in Section 3.4 does not occur if $u>3$, and we can couple $H_{u}(n,p)$ directly to $H_{F}(n,\pi)$ without requiring a hypergraph version of the d-graph $G^{*}$ from Definition 3.3.
@@ 2.1 Induced Edges, Avoidable Configurations and Clean Cycles
When trying to construct a coupling between the random $F$-graph $H={H_{F}}(n,\pi)$ and the copies of $F$ in the random graph $G=G(n,p)$, one invariably encounters the following fundamental problem. Let $H^{\prime}$ be the $F$-graph of $G$, meaning that we include each copy of $F$ in $G$ as an $F$-edge in $H^{\prime}$. For all $0<\pi<1$, $H$ can clearly attain any $F$-graph. But this is not the case for $H^{\prime}$, due to the existence of so-called ’induced $F$-edges’: Assume that there is an $F$-edge $h$, meaning a copy $h$ of $F$, which is not contained in $H$, but that for each (graph) edge $e$ in the copy $h$ of $F$ there exists an $F$-edge $h_{e}$ in $H$ that covers $e$. Then $H^{\prime}$ cannot attain $H$, because if the copies $h_{e}$ of $F$ appear in $G$, then so does the copy $h$ of $F$.
@@ 2.2 Sparse Clean Cycles and Dummy Edges
This leaves us with clean $F$-cycles, which may exist in $H$. Riordan avoids the discussion of these clean $F$-cycles by restricting to the class of graphs $F$ from Theorem 1.3, which he called ‘nice’: A graph with these properties cannot be induced by a clean $F$-cycle. Heckel [13], on the other hand, tweaks Riordan’s coupling to deal with the only type of clean $K_{3}$-cycle which can induce a triangle $K_{3}$, namely the one of length $3$. Since we consider the case of general strictly 1-balanced graphs, we also have to deal with clean $F$-cycles. But as opposed to Heckel, we have to consider all types of clean $F$-cycles. While this is a major complication in its own right, there is one particular type of clean $F$-cycle which is particularly demanding: so-called ‘sparse clean $F$-cycles’, which are given by two $F$-edges, meaning two copies of $F$, that share exactly one (graph) edge (and no other vertices), as in Figure 1.
Figure 1: A sparse clean $F$-cycle, where $F$ is the graph obtained from $K_{4}$ by deleting an edge.
@@ 2.3 Coupling Clean Cycles
With this modification in place, we turn back to the proof in [13], where Heckel couples the clean $F$-cycles in $H$ and their shadow graphs in $G$ so that they match up whp. As opposed to Heckel, we have to couple all types of clean $F$-cycles in $H$, not just one type, with their (selected) shadow graphs in $G^{*}$. For this purpose, as in [13], we have to control the overlaps of pairs of clean $F$-cycles, and the overlaps of their (selected!) shadow graphs in $G^{*}$. Heckel used an exhaustive case distinction and direct computation to gain control. This is not an option for the general case, which brings us to our second main contribution. For the discussion of two overlapping, distinct clean $F$-cycles as well as the discussion of other more complicated $F$-graphs, we observe that these are in fact avoidable configurations. Thus, it is Riordan’s work on avoidable configurations in [24] that immediately gives us control over the overlaps. On the graph side, we establish that shadow graphs $S$ of clean $F$-cycles are strictly balanced, that is, the edge density $e(S)/v(S)$ of any proper subgraph is strictly smaller than the edge density of $S$. Here, it is yet again crucial that we consider the dummy edge when we discuss the shadow graphs of sparse clean $F$-cycles. With this property in place, we easily control the overlaps in the graph case. As with the avoidable configurations above, the fact that these shadow graphs are strictly balanced also helps later on in the proof.
@@ 2.4 The Riordan-Heckel Coupling
To summarize, the coupling in Theorem 1.6 is achieved in the following way:
$c_{F}(G(n,p))=c_{F}(G^{*}(n,p))\overset{\text{Riordan--Heckel}}{\supseteq}H_{F}(n,\pi)$
where we write $c_{F}(G)$ for the set of copies of $F$ contained in $G$. Theorem 1.5 follows from this by observing (e.g. that $H_{F}(n,\pi)$, after merging edges on the same vertex set, is distributed like $H_{r}(n,\pi)$, as outlined in the paragraph above equation (2).
@@ 3.1 Global Constants
3.2 F-graphs
An $F$-graph $H=(V,E)$ is given by the vertex set $V$ and a set $E$ of $F$-edges, that is, copies of $F$ with vertices in $V$. The shadow graph $G=(V,E^{\prime})$ of $H$ is the graph on the same vertices with edge set $E^{\prime}$, which is the union of the edge sets of all $F$-edges in $H$. Formally, let $h=(V_{h},E_{h})$ be the $F$-edges $h\in E$, then we have
@@ 3.2 F-graphs
For a graph $G=(V,E)$, let the $F$-graph $H=(V,E^{\prime})$ of $G$ be defined via the $F$-edge set $E^{\prime}$ of copies of $F$ in $G$. Finally, note that we may think of $r$-uniform hypergraphs as $K_{r}$-graphs and of simple graphs as $K_{2}$-graphs. Thus, it suffices to introduce notions for $F$-graphs (and we take the liberty to allow $F=K_{2}$ for definitions).
@@ 3.3 Induced Edges, Avoidable Configurations And Clean cycles
n(H)=(r1)e(H)+c(H)v(H).n(H)=(r-1)e(H)+c(H)-v(H).
We will tacitly use the following observations on several occasions, to identify minimizers of the nullity in classes of $F$-graphs. Consider two connected $F$-graphs $H_{1}$, $H_{2}$. Let $v$ be a vertex in their vertex overlap such that no $F$-edge incident to $v$ is contained in both $H_{1}$ and $H_{2}$. Obtain the copy $H^{\prime}_{2}$ of $H_{2}$ by substituting $v$ with an entirely new vertex. Then we have $n(H_{1}\cup H^{\prime}_{2})=n(H_{1}\cup H_{2})$ if $v$ is the only vertex in the overlap and $n(H_{1}\cup H^{\prime}_{2})=n(H_{1}\cup H_{2})-1$ otherwise.
Now, let $h$ be an $F$-edge in the overlap of $H_{1}$ and $H_{2}$ and $v$ a vertex in $h$ such that each of the remaining (r1)(r-1) vertices in $h$ is not incident to any other $F$-edge in the overlap. Obtain the copy $H^{\prime}_{2}$ of $H_{2}$ by substituting these (r1)(r-1) vertices with entirely new vertices. Then we have $n(H_{1}\cup H^{\prime}_{2})=n(H_{1}\cup H_{2})$.
Next, let an $F$-edge $h$ be induced by $H$ if it is not contained in $H$, but it is contained in the $F$-graph of the shadow graph of $H$. An $F$-graph $H$ is an avoidable configuration if it is connected and $n(H)\geq 2$. Intuitively, $H$ is an avoidable configuration if it contains at least two cycles. Further, for $k>2$, the $F$-graph $H$ is a clean $F$-cycle of length $k$ if the following holds. There exists an ordering $h_{1},\dots,h_{k}$ of the $F$-edges of $H$ and there are distinct vertices $v_{1},\dots,v_{k}$ such that $h_{i}$ and $h_{i+1}$ overlap exactly in $v_{i}$ for all $i\leq k$ (using $h_{k+1}=h_{1}$), and $h_{i}$, $h_{j}$ do not overlap otherwise. For $k=2$, the $F$-graph $H$ is a clean $F$-cycle of length $2$ if it has exactly two $F$-edges, which overlap in exactly two vertices. The next result from [24] shows that we can control induced $F$-edges by controlling avoidable configurations and clean $F$-cycles. [^1] Note that the statement in [24] is slightly different: the conclusion is only that there exists an avoidable configuration in $H$, or a clean $F$-cycle that induces $E$. However, when following through the proof in [24] it is clear that the avoidable configuration that is found indeed induces $E$.
@@
3.4 Sparse clean cycles and dummy edges
Let $C$ be a clean $F$-cycle. We say that $C$ is sparse if $e(C)=2$ and the two overlap vertices $v_{1},v_{2}$ form an edge $e=\{v_{1},v_{2}\}$ in both $F$-edges $h_{1},h_{2}$ of $C$, that is, the edge $e$ is contained in the copy $h_{1}$ of $F$ as well as in the copy $h_{2}$ of $F$. Otherwise, we say that the clean $F$-cycle $C$ is dense. Thus, we have $e(C^{\prime})=ke(F)$ for the shadow graph $C^{\prime}$ of a dense clean $F$-cycle of length $k$, but $e(C^{\prime})=2e(F)-1$ in the sparse case.
@@ 3.4 Sparse clean cycles and dummy edges
In order to match ${\mathbb{E}}[X]$ and ${\mathbb{E}}[Y]$, we select only a fraction $p$ of the sparse clean $F$-cycles in $G(n,p)$. To model this selection, we introduce the following ‘dummy edges’.
Definition 3.3.
A d-graph $G=(V,U\cup D)$ is given by vertices $V$ and an edge set $E=U\cup D$, where the usual edges $U$ are some subsets of $V$ of size $2$, and the dummy edges $D$ are some sparse clean $F$-cycles with vertices in $V$. The d-graph $G=(V,U\cup D)$ is a clean d-cycle if
  1. 1.
    $(V,U)$ is the shadow graph of a dense clean $F$-cycle and $D=\emptyset$, or
  2. 2.
    $(V,U)$ is the shadow graph of a sparse clean $F$-cycle $C$ and, additionally, $D=\{C\}$, that is, the dummy edge corresponding to $C$ is present.
This definition ensures that for any clean d-cycle $C$ of length $k$ we have $ke(F)$ edges. Finally, the random d-graph $G^{*}(n,p)$ has the vertices $[n]$, and each edge (including dummy edges) is included independently with probability $p$. Note that we obtain $G(n,p)$ from $G^{*}(n,p)$ by discarding the dummy edges.
For d-graphs $G$ we also use $v(G),e(G)=|U\cup D|$, $c(G)$ as before, where $G$ is connected if $G$ without the dummy edges is. We also use the 1-density $d_{1}(G)$ from Equation (1) and the edge density $e(G)/v(G)$, which means that also strictly 1-balanced and strictly balanced d-graphs are defined. The next result is a key argument in the proof.
@@
Now, assume that $G$ is not strictly balanced and let $S\subsetneqq G$ be strictly balanced with maximal density. Assume that $S$ does not cover one of the overlap vertices $v_{i}$. Then we can split at $v_{i}$ to see that $S$ is a subgraph of the shadow graph of a clean $F$-path $P$ with copies $F^{\prime}_{1},\dots,F^{\prime}_{\ell}$ of $F$, that is, $F^{\prime}_{i}$ and $F^{\prime}_{i+1}$ overlap in exactly one vertex for $i<\ell$ and $F^{\prime}_{i}$, $F^{\prime}_{j}$ have no overlap otherwise. Since $S$ is connected, we may assume without loss of generality that $S$ covers all overlap vertices of $P$ (possibly $\ell=1$), and that, using $S_{i}=F^{\prime}_{i}\cap S$, we have $v(S_{i})\geq 2$ for all $i$. It follows that
@@ 4.1 Overlaps
Note that $C_{1}\cup C_{2}$ is an avoidable configuration as it contains two cycles, so Lemma 3.2 completes the bound. ∎
@@
Let $C_{i}$ be the clean $F$-cycle corresponding to $G_{i}$. Further, let $S=G_{1}\cap G_{2}$ be the intersection graph, then we have $\mathbb{E}[X]=\mathcal{O}(n^{f(S)})$ with
@@ 4.2 Coupling
$4\left(\sum_{i}\sum_{j\in B_{i}}\mathbb{E}[X_{i}]\mathbb{E}[X_{j}]+\sum_{i}\sum_{j\in B_{i}\setminus\{i\}}\mathbb{E}[X_{i}X_{j}]\right).$ (4)
@@
Denote by $\mathscr{C}$ the set of all potential copies of clean $F$-cycles of length at most $e(F)$ in $H_{F}(n,\pi)$. Each $C\in\mathscr{C}$ corresponds bijectively to a potential clean d-cycle in $G^{*}(n,p)$, and thus we will use $\mathscr{C}$ as an index set for both clean $F$-cycles and clean d-cycles by a slight abuse of notation.
For $C\in\mathscr{C}$ let $X_{C}$ be the indicator for the event that $C$ is in ${H_{F}}(n,\pi)$. Furthermore, let $B_{C}$ be the set of all clean $F$-cycles that are not vertex disjoint from $C$. Note that for $C^{\prime}\in B_{C}$ we have $\mathbb{E}[X_{C}X_{C^{\prime}}]\geq\mathbb{E}[X_{C}]\mathbb{E}[X_{C^{\prime}}]$, thus the term in (4) is upper bounded by
$8\sum_{C\in\mathscr{C}}\ \sum_{C^{\prime}\in B_{C}\setminus\{C\}}\mathbb{E}[X_{C}X_{C^{\prime}}]+4\sum_{C\in\mathscr{C}}\mathbb{E}[X_{C}]^{2}.$
This is $o(1)$: Indeed, collecting terms in the double-sum according to the isomorphism class of $C\cup C^{\prime}$ (of which there are $\mathcal{O}(1)$ many) yields a bound of $\mathcal{O}(n^{-1+2e(F)\varepsilon})=o(1)$ by Lemma 4.1. Similarly, splitting up the second sum according to the isomorphism class of $C$ yields $\mathcal{O}\big{(}n^{2e(F)(\varepsilon-2(r-1)}\big{)}=o(1)$, analagous to the proof of Lemma 3.2. Hence there is a coupling of of $(X_{C})_{C\in\mathscr{C}}$ and independent Poisson variables $(Y_{C})_{C\in\mathscr{C}}$ with parameters $\mathbb{E}[X_{C}]$ such that $(X_{C})_{C}=(Y_{C})_{C}$ whp.
Relying on Lemma 4.2, we proceed analogously for $G^{*}(n,p)$ to arrive at a coupling of the indicators $(X^{\prime}_{C})_{C}$ and independent Poisson variables $(Y^{\prime}_{C})_{C}$ with parameters $\mathbb{E}[X^{\prime}_{C}]$ such that $(X^{\prime}_{C})_{C}=(Y^{\prime}_{C})_{C}$ whp.
Now, note that $\mathbb{E}[X_{C}]=(1-n^{-\delta})^{e(C)}\mathbb{E}[X^{\prime}_{C}]$. Therefore,
$d_{TV}\big{(}(Y_{C})_{C},(Y^{\prime}_{C})_{C}\big{)}\leq\sum_{C\in\mathscr{C}}d_{TV}(Y_{C},Y^{\prime}_{C})\leq\sum_{C\in\mathscr{C}}\left((1-n^{-\delta})^{-e(C)}-1\right)\mathbb{E}[X_{C}]=\mathcal{O}\big{(}n^{-\delta+\varepsilon e(F)}\big{)}$
by Lemma 3.2 after collecting summands with isomorphic $C$ and noting that $n(C)=1$. Thus for $\varepsilon<\delta/e(F)$, we can couple $(Y_{C})_{C}$ with $(Y^{\prime}_{C})_{C}$ such that $(Y_{C})_{C}=(Y^{\prime}_{C})_{C}$ whp. ∎